OpenStudy (graciebugg):

Will Fan and Medal..... Might even Add a Testimonial. Which shows the expression below in simplified form? (3 × 10^8) ÷ (5 × 10^4) A.6 × 10^5 B.6 × 10^4 C.-2 × 10^3 D. 6 × 10^3

OpenStudy (graciebugg):

@lochana

OpenStudy (graciebugg):

I say C.

OpenStudy (lochana):

\[\frac{3.10^8}{5.10^4} = \frac{3}{5}10^4\]

OpenStudy (lochana):

no. that's not the answer

OpenStudy (lochana):

\[\frac{3}{5}10^4 = \frac{6}{10}10^4\]

OpenStudy (lochana):

\[6.10^3\]

OpenStudy (graciebugg):

Ok... I only said that because of negative 2... I suck at math

OpenStudy (lochana):

ah that's fine. try more problems . you will be better:)

OpenStudy (graciebugg):

So it's D?

OpenStudy (lochana):

yes